OpenStudy (anonymous):

Can someone help me solve this equations graphically? 3y+2x=4 4y+5x=3

OpenStudy (phi):

the first step is graph the two equations 3y+2x=4 I would look for 2 points on this line. pick a y, for example y=0. now find x. can you do that?

OpenStudy (anonymous):

Find x for both of the equations?

OpenStudy (phi):

no. one equation at a time. first let's work on 3y+2x=4

OpenStudy (anonymous):

Ok so 3(0)+2x=4 so x+2, right?

OpenStudy (phi):

you mean x=2. so one point on this line (it's a line by the way) is (2,0) now try y=2 (I think we get a nicer number than if we used y=1)

OpenStudy (anonymous):

Yes! Sorry. so x=-4!

OpenStudy (anonymous):

(2,-4)

OpenStudy (anonymous):

(-4,2)

OpenStudy (phi):

I got something different. how did you get -4?

OpenStudy (phi):

3y+2x=4 with y=2 3*2+2x=4 6+2x=4 2x= 4-6 2x=-2 x=-1

OpenStudy (anonymous):

Oh, my bad. Thanks though, So how many points do you think i should make for the graph?

OpenStudy (phi):

you only need 2 points.

OpenStudy (anonymous):

Ah, true. Thanks! I really appreciated your help. :)

OpenStudy (phi):

so the 2 points you have for the first equation are (2,0) and (-1,2)

OpenStudy (phi):

you plot those points and draw a line through them. now we need 2 points on the other line 4y+5x=3 another way to plot this is put it in slope intercept form y= mx+b solve for y: 4y= -5x+3 y= -5/4 x +3/4 its y-intercept is 3/4 so one point is (0, 3/4)

OpenStudy (phi):

using y= -5/4 x +3/4 with x=1 we get y= -5/4 +3/4 = -2/4 = -1/2 so (1, -1/2) is another point two points for the 2nd line: (0, 3/4) (1, -1/2)

OpenStudy (phi):

can you plot those points?

OpenStudy (phi):

if you can, then plot the points, draw a line through them. Look for where the 2 lines cross. that is the "solution" I would read the (x,y) point where the lines meet, and test that they work in both equations (as a check)